OpenStudy (anonymous):

Solve. 1.7x + 8 − 1.62x = 0.4 x − 0.32 + 8

OpenStudy (anonymous):

0.32x = 0.32 x = 1

OpenStudy (anonymous):

SHO WORK JIMMY

OpenStudy (anonymous):

0.4x - 1.7x + 1.62x = 8 - 8 + 0.32 0.32x = 0.32

OpenStudy (anonymous):

ok?